PHY 203 Introduction to Physics NAME:
(5.1) A particle has constant acceleration a0 = −4 m/s
2. It is at rest at x0 = 128 m at
t = 0 s.
(a) (i) Determine the velocity of the particle at t = 2 s.
(ii) Determine the position of the particle at t = 2 s.
(b) (i) Determine the velocity of the particle at t = 4 s.
(ii) Determine the position of the particle at t = 4 s.
(c) (i) Determine the velocity of the particle at t = 8 s.
(ii) Determine the position of the particle at t = 8 s.
(5.2) A particle has constant acceleration a0 = −4 m/s
2. It is at rest at x0 = 128 m at
t = 0 s.
(a) Determine the speed of the particle just as it passes the point x = 120 m.
(b) Determine the speed of the particle just as it passes the point x = 110 m.
(c) Determine the speed of the particle just as it passes the point x = 56 m.
(d) Determine the speed of the particle just as it passes the point x = 30 m.
Answer #1

Yo = 126m , at t os q 0 o= - 4 m/s2 â . U = Uxtat Un = o-4.0x2 = -8.0 m/s cis X(2) = xo + 4 axt2 = 128 - 4x4.0x12,2 = 128-8 =